OpenStudy (anonymous):

find the perimeter of a rectangular object with the length of (square root 75) feet and the width (square root 48 feet)???????????????

OpenStudy (anonymous):

perimeter is the total length of all the edges of the object so P=2*sqrt(75)+2*sqrt(48) P=2*sqrt(3*25)+2*sqrt(3*16) P=2*5sqrt(3)+2*4sqrt(3) P=10sqrt(3)+8sqrt(3) P=18sqrt(3)

OpenStudy (anonymous):

is that how to do it or the answer?

OpenStudy (anonymous):

that's how to do it. but if you struggle with changing sqrt(75) ans sqrt48 into simpler form then I can get you more detail

OpenStudy (anonymous):

ok, so i think you have problem with that. sqrt = square root sqrt(75)=sqrt(3*5*5)=sqrt(3*5^2)=sqrt(3)*sqrt(5^2)=5*sqrt(3) 48 is similar

OpenStudy (anonymous):

i dont get what the peremiter would be tho of the two ?

OpenStudy (anonymous):

rectangular has 2 longer edges and 2 shorter edges Perimeter = total lengths of 2 longer edges + total length of 2 shorter edges you know the longer edge is sqrt75 ft and shorter edge is sqrt48 ft, so parameter: P= sqrt75 + sqrt75 + sqrt48 +sqrt48 = 2sqrt75 + 2sqrt48

OpenStudy (anonymous):

rectangle lol

OpenStudy (anonymous):

so it would be 5 sqrt 3 plus 5 sqrt 3 plus 4 sqrt 3 or something

OpenStudy (anonymous):

5 sqrt 3 plus 5 sqrt 3 plus 4 sqrt 3 plus 4 sqrt 3 since 2 edges each.

OpenStudy (anonymous):

do i add them together for the answer

OpenStudy (anonymous):

yeah, you will get 18 sqrt3

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

no problem, as long as you understand how it works, there shouldn't be any other problems

OpenStudy (anonymous):

nope do you no what a house would look like if using those measures

OpenStudy (anonymous):

what do you mean? the base should be a nice squared rectangle